# From ABC to Effective Roth and Ridout Constants for Cubic Roots

### v1

@@ 2.14 Bounds for the approximation gain of ABC hits from resulting equations
approximation gain = $\frac{ln({p_{n}}^{m})}{ln(-d_{n}q_{n}kp_{n})}\leq$ $\frac{m}{2}$
@@
So the ABC gains can be seperated for resulting equation in approximation gains and power gains. The approximation gain can be bounded by $\frac{3}{2}$ for 3rd roots. To prove a bound for the quality the power gains $\frac{ln({-p_{n}q_{n}d_{n}k})}{ln(rad(-d_{n}q_{n}kp_{n}))}$ must be bounded as well, which seems to be very difficult even for special cases like $\sqrt[3]{2}$ (remarkable are here the second equation 128 = 125 + 3 with a power gain of 1,4 and an approximation gain of 1,01 and the 5th equation with a power gain of 1,37 and an approximation gain of 0,99).
@@
$c<L_{\varepsilon}\cdot rad(abc)^{3+\varepsilon}.$
@@
power gain = $\frac{ln({-p_{n}q_{n}d_{n}k})}{ln(rad(-d_{n}q_{n}kp_{n}))}$
$\leq\frac{ln({k}^{2}{p_{n}}^{3})}{ln(rad(-d_{n}q_{n}kp_{n}))}$
@@
= $\frac{ln({k}^{2}{p_{n}}^{3})}{ln((rad(-d_{n}q_{n}kp_{n})^{3+\varepsilon})^{\frac{1}{3+\varepsilon}})}$
@@
One concept to prove the ABC conjecture might then be to prove that both gains are bounded seperately. We conjecture that the approximation gain is bounded by $\frac{3}{2}$.

### v2

@@ 2.16 General case if the formula can be used
approximation gain = $\frac{ln({p_{n}}^{m})}{ln(d_{n}q_{n}kp_{n})}\leq$ $\frac{m}{2}$
@@
So the ABC gains can be seperated for resulting equation in approximation gains and power gains. The approximation gain can be bounded by $\frac{3}{2}$ for 3rd roots. To prove a bound for the quality the power gains $\frac{ln({p_{n}q_{n}d_{n}k})}{ln(rad(d_{n}q_{n}kp_{n}))}$ must be bounded as well, which seems to be very difficult even for special cases like $\sqrt[3]{2}$ (remarkable are here the second equation 128 = 125 + 3 with a power gain of 1.4 and an approximation gain of 1.01 and the 5th equation with a power gain of 1.37 and an approximation gain of 0.99).
@@
$\Large c<L_{\varepsilon}\cdot rad(abc)^{3+\varepsilon}.$
@@
power gain = $\frac{ln({p_{n}q_{n}d_{n}k})}{ln(rad(d_{n}q_{n}kp_{n}))}$
$\leq\frac{ln({k}^{2}{p_{n}}^{3})}{ln(rad(d_{n}q_{n}kp_{n}))}$
@@
= $\frac{ln({k}^{2}{p_{n}}^{3})}{ln((rad(d_{n}q_{n}kp_{n})^{3+\varepsilon})^{\frac{1}{3+\varepsilon}})}$
@@
One concept to prove the ABC conjecture might then be to prove that both gains are bounded seperately. We conjecture that the approximation gain is bounded by $\frac{3}{2}$ and the power gain by 3.